# Topologically Protected Surface Altermagnetism on Antiferromagnets

### v1

@@ Abstract
Symmetry requirements alone do not guarantee robust surface AM, since the surface state lacks topological protection and so can be destroyed with a symmetry-respecting perturbation. We proposed two different mechanisms for topological surface AM: (i) AM drumhead states in a Dirac nodal line semimetal and (ii) AM Fermi arc states in a Dirac semimetal. The bulk-boundary correspondence then guarantees robust surface AM.
@@
The spin-splitting of the altermagnetic surface states cannot be identified via global properties like the band structure because opposite surface states compensate one another. Rather, the effect is only visible in the surface spin spectral density, which can be resolved by surface-focused, spin- and angle resolved photoemission spectroscopy [24, 25, 29]. Alternatively, we suggest that surface and spin-resolved transport measurements should be able to capture the local spin splitting [2, 12, 18].
@@
Several existing materials fulfill the symmetry requirements to exhibit surface AM, and just need to be analyzed at the surface. In CuMnAs ab-initio calculation revealed a spin-split Fermi arc state under strain, in the related compound CuMnP it is yet to be confirmed [36]. Furthermore, $C_{2}\parallel\bm{t}_{z}]-$protected AFMs consisting of spin-inverted AM layers [38], e.g. induced by orbital ordering [31, 38, 8], should show surface AM. In general, studies have focused on inducing AM in materials by explicitly break inversion symmetry, e.g. through an electric field [37]. As evidenced here, these materials are worth reanalyzing by considering the local surface physics.
@@
Our work substantially enlarges the material landscape for AM and identifies topology as a resource for engineering robust spin-polarized electronic states in spin-degenerate AFMs. We therefore expect topological surface AM to provide a versatile route towards robust antiferromagnetic spintronics and to be a fertile ground for unconventional surface-driven phenomena.
@@
Data and code availability—Code and data related to this paper are available from the authors upon reasonable request. This work used the python package blochK [32].
@@
Appendix A Spin conductance and spin splitter effect
In our work we study a layer-resolved spin splitter effect, which is based on the the layer-resolved (spin) conductivity. Our calculation is based on the following setup: An uniform electric field is applied on the entire slab (parallel to the surface), the total current of each pair of layers (one unit cell) is measured.
@@ Appendix A Spin conductance and spin splitter effect
$\tilde{\sigma}_{\mu\nu}^{\alpha}(z)=\frac{e}{\pi V}\mathrm{Re}\sum_{\bm{k},m,n}\frac{\langle u_{n}(\bm{k})|J^{\alpha}_{\mu}(z)|u_{m}(\bm{k})\rangle\langle u_{m}(\bm{k})|v_{\nu}|u_{n}(\bm{k})\rangle\Gamma^{2}}{\left(\epsilon_{n}(\bm{k})^{2}+\Gamma^{2}\right)\left(\epsilon_{m}(\bm{k})^{2}+\Gamma^{2}\right)}$ (6)
@@
$\sigma_{\mu\nu}^{\uparrow}(z)$ $=\left(\tilde{\sigma}^{0}_{\mu\nu}(z)+\left[\tilde{\sigma}^{z}_{\mu\nu}(z)+\tilde{\sigma}^{z}_{\nu\mu}(z)\right]/2\right)/2$ (9)
$\sigma_{\mu\nu}^{\downarrow}(z)$ $=\left(\tilde{\sigma}^{0}_{\mu\nu}(z)-\left[\tilde{\sigma}^{z}_{\mu\nu}(z)+\tilde{\sigma}^{z}_{\nu\mu}(z)\right]/2\right)/2.$ (10)
@@
$\alpha_{x}=2\arctan\left(\frac{\sigma_{xy}^{\uparrow}-\sigma_{xy}^{\downarrow}}{\sigma_{xx}^{\uparrow}+\sigma_{xx}^{\downarrow}}\right)$ (11)
$\alpha_{y}=2\arctan\left(\frac{\sigma_{xy}^{\uparrow}-\sigma_{xy}^{\downarrow}}{\sigma_{yy}^{\uparrow}+\sigma_{yy}^{\downarrow}}\right)$ (12)
@@
$\alpha=2\arctan\left(\frac{\sigma_{xx}^{\uparrow}-\sigma_{xx}^{\downarrow}}{\sigma_{xx}^{\uparrow}+\sigma_{xx}^{\downarrow}}\right).$ (13)
Appendix B Emergence of Spontaneous AM in a Hubbard Model
@@
$H_{U}=$ $-t\sum_{\langle ij\rangle,s}c_{i,s}^{\dagger}c_{j,s}-t^{\prime}\sum_{\langle\langle ij\rangle\rangle\perp\hat{e}_{z},s}(1+(-1)^{i}\delta)c_{i,s}^{\dagger}c_{j,s}+U\sum_{i}c_{i,\uparrow}^{\dagger}c_{i,\uparrow}c_{i,\downarrow}^{\dagger}c_{i,\downarrow}$ (14)
whose hopping structure is identical to Eq. (1). For dominant NN exchange, $t>t^{\prime},\delta$, around half-filling it is well understood that superexchange leads to an AFM state where nearest neighbor sites have opposite magnetization $\tilde{m}$. This motivates the mean-field ansatz
$\langle c_{i,s}^{\dagger}c_{i,s}\rangle=n/8+(-1)^{i+s}\tilde{m}$. (15)
The resulting magnetic state is an $(0,0,\pi)$-AFM, because there are already 2 sublattices in the $k_{x}$-$k_{y}$-plane due to the NNN hopping anisotropy $\delta$. Hence, the magnetic unit cell consists of 4 sites: 2 layers (subspace $\tau$) times 2 identical sublattices (subspace $\sigma$) per layer.
B.1 MF decoupling in 3D
Mean field decoupling the Hubbard interaction using Eq. (15) and $AB\rightarrow\langle A\rangle B+A\langle B\rangle-\langle A\rangle\langle B\rangle$ gives
$U\sum_{i}c_{i,\uparrow}^{\dagger}c_{i,\uparrow}c_{i,\downarrow}^{\dagger}c_{i,\downarrow}\rightarrow U\sum_{i,s}c_{i,s}^{\dagger}c_{i,s}\left(\frac{n}{8}-(-1)^{i+s}\tilde{m}\right)+4UN\left(\frac{n^{2}}{64}-\tilde{m}^{2}\right).$ (16)
@@ B.1 MF decoupling in 3D
$-2\delta\sigma_{z}\left[\cos(k_{x}+k_{y})-\cos(k_{x}-k_{y})\right]-\sigma_{z}s_{z}\tau_{z}m$ (17)
@@
$\pm\left.2\left[(m^{2}+4t^{2}\cos^{2}k_{z})(4t^{2}\left(\cos k_{x}+\cos k_{y}\right)^{2}+4\delta^{2}\left(\cos(k_{x}+k_{y})-\cos(k_{x}-k_{y})\right)^{2})\right]^{1/2}\right]^{1/2}.$ (18)
The spin-degeneracy is a result of the combined time reversal inversion symmetry $\mathcal{T}\parallel\mathcal{I}]$. In our model time reversal is ${\mathrm{i}}s_{y}\mathcal{K}$ ($\mathcal{K}$ is complex conjugation) with $\bm{k}\rightarrow-\bm{k}$ and inversion is $\tau_{x}$ with $\bm{k}\rightarrow-\bm{k}$. Hence, the invariance can be explicitly observed by the fact that $h(\bm{k})$ is real and commutes with $s_{y}\tau_{x}$.
B.2 MF decoupling of the slab
Now we consider a slab, i.e. a finite amount of layers in $z$-direction. In the main manuscript we considered a constant density and magnetization over all layers. Here, we show that there is a layer dependence of density and magnetization, which does however not affect our conclusions.
@@ B.2 MF decoupling of the slab
$\langle c_{(x,y,z),s}^{\dagger}c_{(x,y,z),s}\rangle=n_{z}/8+(-1)^{x+y+z+s}\tilde{m}_{z}/2.$ (19)
@@
$U\sum_{i}c_{i,\uparrow}^{\dagger}c_{i,\uparrow}c_{i,\downarrow}^{\dagger}c_{i,\downarrow}\rightarrow U\sum_{k_{x},k_{y},z,s}\left(\frac{n_{z}}{8}\bm{\Psi}_{k_{x},k_{y},z}^{\dagger}\cdot\bm{\Psi}_{k_{x},k_{y},z}-(-1)^{z}\frac{\tilde{m}}{2}\bm{\Psi}_{k_{x},k_{y},z}^{\dagger}\tau_{z}\sigma_{z}\bm{\Psi}_{k_{x},k_{y},z}\right)-UL_{x}L_{y}\sum_{z}\left(\frac{n_{z}^{2}}{16}-\tilde{m}_{z}^{2}\right).$ (20)
@@
$H=UL_{x}L_{y}\sum_{z}\left(\tilde{m}_{z}^{2}-\frac{n_{z}^{2}}{16}\right)+\sum_{k_{x},k_{y},z}\bm{\Psi}_{k_{x},k_{y},z}^{\dagger}h(k_{x},k_{y},z)\bm{\Psi}_{k_{x},k_{y},z}-t\bm{\Psi}_{k_{x},k_{y},z}^{\dagger}\cdot\bm{\Psi}_{k_{x},k_{y},z+1}+\text{h.c.}$ (21)
@@
$-2\delta\sigma_{z}\left[\cos(k_{x}+k_{y})-\cos(k_{x}-k_{y})\right]+\frac{U}{8}n_{z}-\sigma_{z}s_{z}(-1)^{z}U\tilde{m}_{z}$ (22)
@@
We performed combined restricted (in $x$ and $y$ direction) and unrestricted (in $z$ direction) Hartree–Fock mean field calculations. We compared the energy of the solution of the ansatz above to non-magnetic and ferromagnetic solutions. For sufficiently large $U$ and close to half-filling the $(0,0,\pi)$ AFM with a weak $z$-dependence, see Fig. 4, is consistently the lowest energetic state.
@@
Figure 4: Real space mean field simulations of the slab confirm that assuming constant mean fields is a valid assumption. Mean field result of Eq. (21) for $t^{\prime}=0.3t$, $U=6t$, $\delta=1$, around half-filling $n=0.54$. (a) The AFM magnetization $m_{z}$ tends to increase slightly towards the surface, whereas the density $n_{z}$ is almost constant. (b) Comparison of the spin splitter effects for layer-dependent mean field results (solid) and using the average values for the mean fields (dashed). (c-f) Comparison of the surface (c-d) and bulk (e-f) spin spectral density (opacity encodes absolute spectral density and the red to blue encodes the spin character) when using the layer-dependent mean field values (c,e) and their mean values (d,f).
Appendix C Minimal model for orthorhombic CuMnAs
@@
The bulk polymorph of CuMnAs has orthorhombic space group Pnma (SG #62). Note a tetragonal structure with space group P4/nmm (SG #129) also exists, but is not the focus of this work [50]. Pnma is a non-symmorphic space group generated by three mutually orthogonal mirror symmetries $M_{a}\{\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2}\}$, $M_{b}\{0,\tfrac{1}{2},0\}$ and $M_{c}\{\tfrac{1}{2},0,\tfrac{1}{2}\}$. We take the coordinate system such that $(y,z,x)$ correspond to $(a,b,c)$ axes. All three atoms (Cu,Mn,As) sit on 4c positions [13]. Magnetic moments are located at the Mn atoms in an AFM pattern which preserves $\mathcal{IT}$ symmetry. In addition, in the absence of SOC the symmetry of the magnetic order can be analyzed separately in spin and spatial sectors, and the magnetic order is odd under the mirror $M_{c}$ but even under $M_{a}$ and $M_{b}$. At the surface normal to the $b$ axis, $\mathcal{IT}$ is broken but the combined symmetry $M_{c}\mathcal{T}$ is preserved in the absence of SOC, enforcing a zero total moment [51] and thus qualifying as an altermagnetic surface.
@@
Figure 5: Tight-binding model for Wyckoff position 4a of Pmna. Filled (empty circles) are sites on even (odd) layers accounted by $\tau$-Pauli matrices, yellow (purple) circles are sites of the A (B) sublattices accounted by $\sigma$-Pauli matrices (a) In-plane and out-of-plane hoppings which preserve all three accidental translations in Eqs. (28)-(30). (b) Further neighbor hoppings which keep only $\bm{t}_{a/2,b/2,c/2}$. (c) Further neighbor hoppings which keep only $\bm{t}_{a/2,c/2}$. (d) (110) hopping $t^{\prime}$, which gaps out the nodal line into 4 Dirac cones. (e-h) Spin-dependent hoppings corresponding to the magnetic order with odd $M_{c}\{\tfrac{1}{2},0,\tfrac{1}{2}\}$ symmetry.
The nodal line occurs at $k_{z}=0$, and is protected by the mirror $M_{b}$ which leaves the $k_{z}=0$ plane invariant. The nodal line occurs within a subset of four bands, which come in pairs of opposite $M_{b}$ eigenvalues. Since Mn sits in Wyckoff position 4c, which is invariant under $M_{b}$, a model with a single orbital at 4c necessarily has all four bands which are even or odd under $M_{b}$, which means it is impossible to build a four band model for the nodal line. This means the Wannier centre for such set of four bands must be in another position with multiplicity 4, which can be 4a or 4b. Both positions are not invariant under $M_{b}$ and either can be picked, so we choose 4a. The model is layered in the $b$ direction, with orbitals at $(0,0,0)$, $(\tfrac{1}{2},\tfrac{1}{2},0)$ in the first layer and $(0,0,\tfrac{1}{2})$, $(\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2})$ in the second layer. These sites are shown in Fig. 5(a) and labelled (1-4) in this order. Taking the coordinate system where $(a,b,c)\leftrightarrow(k_{y},k_{z},k_{x})$ and in the basis $(c_{1},c_{2},c_{3},c_{4})$ corresponding to the four sites in Fig. 5(a), we define sublattice $\sigma_{i}$, layer $\tau_{i}$, and spin $s_{i}$ Pauli matrices.
$M_{a}:$ $H(k_{x},k_{y},k_{z})$ $=\sigma_{x}\tau_{x}H(k_{x},-k_{y},k_{z})\sigma_{x}\tau_{x}$ (23)
$M_{b}:$ $H(k_{x},k_{y},k_{z})$ $=\tau_{x}H(k_{x},k_{y},-k_{z})\tau_{x}$ (24)
$M_{c}:$ $H(k_{x},k_{y},k_{z})$ $=\sigma_{x}H(-k_{x},k_{y},k_{z})\sigma_{x}$ (25)
$\mathcal{I}:$ $H(\bm{k})$ $=H(-\bm{k})$ (26)

### v2

@@ Abstract
Symmetry requirements alone do not guarantee robust surface AM, because without topological protection surface states can be destroyed by a symmetry-respecting perturbation. We proposed two different mechanisms for robust topological surface AM: (i) AM drumhead states in a Dirac nodal line semimetal and (ii) AM Fermi arc states in a Dirac semimetal. The bulk-boundary correspondence then guarantees robust surface AM.
@@
The spin-splitting of altermagnetic surface states cannot be identified via global electronic or transport properties because opposite surface states compensate one another. Rather, the effect is only visible in the surface spin spectral density, which can be resolved by surface-focused, spin- and angle resolved photoemission spectroscopy [26, 27, 31]. Alternatively, we suggest that surface and spin-resolved transport measurements should be able to capture the local spin splitting [2, 13, 20].
@@
Several existing materials fulfill the symmetry requirements to exhibit surface AM, and just need to be analyzed at the surface. In CuMnAs ab-initio calculation revealed a spin-split Fermi arc state under strain, in the related compound CuMnP it is yet to be confirmed [38]. Furthermore, $C_{2}\parallel\bm{t}_{z}]-$protected AFMs consisting of spin-inverted AM layers [40], e.g. induced by orbital ordering [33, 40, 9], should show surface AM. In general, materials where studies have focused on inducing AM by explicitly break inversion symmetry, e.g. through an electric field [39, 18] are worth reanalyzing by considering the local surface physics.
@@
Our work substantially enlarges the material landscape for AM and identifies topology as a resource for engineering robust spin-polarized electronic states at the surface of spin-degenerate AFMs. This is especially pertinent since a majority of magnetic materials are topological [elcoro_magnetic_2021], and AFMs are the most common form of magnetic ordering [nmec_antiferromagnets_2018]. We therefore expect topological surface AM to provide a provide a new avenue of spintronics applications in AFMs and to provide a fertile ground for unconventional surface-driven phenomena.
@@
Data and code availability—All code and data related to this paper, including the python package blochK [34], are publicly available [alter_surf].
@@
Supplemental Material: Topologically Protected Surface Altermagnetism on Antiferromagnets
I Spin conductance and spin splitter effect
In our work we study a layer-resolved spin splitter effect, which is based on the the layer-resolved (spin) conductivity. Our calculation is based on the following setup: An uniform electric field is applied on the entire slab (parallel to the surface), the spin-resolved current of each pair of layers (one bulk unit cell) is measured.
@@ I Spin conductance and spin splitter effect
$\tilde{\sigma}_{\mu\nu}^{\alpha}(z)=\frac{e}{\pi V}\mathrm{Re}\sum_{\bm{k},m,n}\frac{\langle u_{n}(\bm{k})|J^{\alpha}_{\mu}(z)|u_{m}(\bm{k})\rangle\langle u_{m}(\bm{k})|v_{\nu}|u_{n}(\bm{k})\rangle\Gamma^{2}}{\left(\epsilon_{n}(\bm{k})^{2}+\Gamma^{2}\right)\left(\epsilon_{m}(\bm{k})^{2}+\Gamma^{2}\right)}$ (S1)
@@
$\sigma_{\mu\nu}^{\uparrow}(z)$ $=\left(\tilde{\sigma}^{0}_{\mu\nu}(z)+\left[\tilde{\sigma}^{z}_{\mu\nu}(z)+\tilde{\sigma}^{z}_{\nu\mu}(z)\right]/2\right)/2$ (S4)
$\sigma_{\mu\nu}^{\downarrow}(z)$ $=\left(\tilde{\sigma}^{0}_{\mu\nu}(z)-\left[\tilde{\sigma}^{z}_{\mu\nu}(z)+\tilde{\sigma}^{z}_{\nu\mu}(z)\right]/2\right)/2.$ (S5)
@@
$\alpha_{x}=2\arctan\left(\frac{\sigma_{xy}^{\uparrow}-\sigma_{xy}^{\downarrow}}{\sigma_{xx}^{\uparrow}+\sigma_{xx}^{\downarrow}}\right)$ (S6)
$\alpha_{y}=2\arctan\left(\frac{\sigma_{xy}^{\uparrow}-\sigma_{xy}^{\downarrow}}{\sigma_{yy}^{\uparrow}+\sigma_{yy}^{\downarrow}}\right)$ (S7)
@@
$\alpha=2\arctan\left(\frac{\sigma_{xx}^{\uparrow}-\sigma_{xx}^{\downarrow}}{\sigma_{xx}^{\uparrow}+\sigma_{xx}^{\downarrow}}\right).$ (S8)
II Emergence of Spontaneous AM in a Hubbard Model
@@
$H_{U}=$ $-t\sum_{\langle ij\rangle,s}c_{i,s}^{\dagger}c_{j,s}-t^{\prime}\sum_{\langle\langle ij\rangle\rangle\perp\hat{e}_{z},s}(1+(-1)^{i}\delta)c_{i,s}^{\dagger}c_{j,s}+U\sum_{i}c_{i,\uparrow}^{\dagger}c_{i,\uparrow}c_{i,\downarrow}^{\dagger}c_{i,\downarrow}$ (S9)
whose hopping structure is identical to Eq. (1). For dominant NN exchange, $t>t^{\prime},\delta$ and around half-filling it is well understood that superexchange leads to an AFM state where nearest neighbor sites have opposite magnetization $\tilde{m}$. This motivates the mean-field ansatz
$\langle c_{i,s}^{\dagger}c_{i,s}\rangle=n/8+(-1)^{i+s}\tilde{m}$. (S10)
The resulting magnetic state is a $(0,0,\pi)$-AFM, because there are already 2 sublattices in the $k_{x}$-$k_{y}$-plane due to the NNN hopping anisotropy $\delta$. Hence, the magnetic unit cell consists of 4 sites: 2 crystallographically identical layers (subspace $\tau$) times 2 sublattices (subspace $\sigma$) per layer.
II.1 MF decoupling in 3D
Mean field decoupling the Hubbard interaction using Eq. (S10) and $AB\rightarrow\langle A\rangle B+A\langle B\rangle-\langle A\rangle\langle B\rangle$ gives
$U\sum_{i}c_{i,\uparrow}^{\dagger}c_{i,\uparrow}c_{i,\downarrow}^{\dagger}c_{i,\downarrow}\rightarrow U\sum_{i,s}c_{i,s}^{\dagger}c_{i,s}\left(\frac{n}{8}-(-1)^{i+s}\tilde{m}\right)+4UN\left(\frac{n^{2}}{64}-\tilde{m}^{2}\right).$ (S11)
@@ II.1 MF decoupling in 3D
$-2\delta\sigma_{z}\left[\cos(k_{x}+k_{y})-\cos(k_{x}-k_{y})\right]-\sigma_{z}s_{z}\tau_{z}m$ (S12)
@@
$\pm\left.2\left[(m^{2}+4t^{2}\cos^{2}k_{z})(4t^{2}\left(\cos k_{x}+\cos k_{y}\right)^{2}+4\delta^{2}\left(\cos(k_{x}+k_{y})-\cos(k_{x}-k_{y})\right)^{2})\right]^{1/2}\right]^{1/2}.$ (S13)
The spin-degeneracy is a result of the combined time reversal inversion symmetry $\mathcal{TI}$. In our model time reversal is ${\mathrm{i}}s_{y}\mathcal{K}$ ($\mathcal{K}$ is complex conjugation) with $\bm{k}\rightarrow-\bm{k}$ and inversion is $\tau_{x}$ with $\bm{k}\rightarrow-\bm{k}$. Hence, the invariance can be explicitly observed by the fact that $h(\bm{k})$ is real and commutes with $s_{y}\tau_{x}$.
II.2 MF decoupling of the slab
Now we consider a slab, i.e., a finite amount of layers in $z$-direction. In the main manuscript we considered a constant density and magnetization over all layers. Here, we show that there is a layer dependence of density and magnetization, which does not affect our conclusions.
@@ II.2 MF decoupling of the slab
$\langle c_{(x,y,z),s}^{\dagger}c_{(x,y,z),s}\rangle=n_{z}/8+(-1)^{x+y+z+s}\tilde{m}_{z}/2.$ (S14)
@@
$U\sum_{i}c_{i,\uparrow}^{\dagger}c_{i,\uparrow}c_{i,\downarrow}^{\dagger}c_{i,\downarrow}\rightarrow U\sum_{k_{x},k_{y},z,s}\left(\frac{n_{z}}{8}\bm{\Psi}_{k_{x},k_{y},z}^{\dagger}\cdot\bm{\Psi}_{k_{x},k_{y},z}-(-1)^{z}\frac{\tilde{m}}{2}\bm{\Psi}_{k_{x},k_{y},z}^{\dagger}\tau_{z}\sigma_{z}\bm{\Psi}_{k_{x},k_{y},z}\right)-UL_{x}L_{y}\sum_{z}\left(\frac{n_{z}^{2}}{16}-\tilde{m}_{z}^{2}\right).$ (S15)
@@
$H=UL_{x}L_{y}\sum_{z}\left(\tilde{m}_{z}^{2}-\frac{n_{z}^{2}}{16}\right)+\sum_{k_{x},k_{y},z}\bm{\Psi}_{k_{x},k_{y},z}^{\dagger}h(k_{x},k_{y},z)\bm{\Psi}_{k_{x},k_{y},z}-t\bm{\Psi}_{k_{x},k_{y},z}^{\dagger}\cdot\bm{\Psi}_{k_{x},k_{y},z+1}+\text{h.c.}$ (S16)
@@
$-2\delta\sigma_{z}\left[\cos(k_{x}+k_{y})-\cos(k_{x}-k_{y})\right]+\frac{U}{8}n_{z}-\sigma_{z}s_{z}(-1)^{z}U\tilde{m}_{z}$ (S17)
@@
We performed combined restricted (in $x$ and $y$ direction) and unrestricted (in $z$ direction) Hartree–Fock mean field calculations. We compared the energy of the solution of the ansatz above to non-magnetic and ferromagnetic solutions. For sufficiently large $U$ and close to half-filling the $(0,0,\pi)$ AFM with a weak $z$-dependence, see Fig. S1, is consistently the lowest energetic state.
@@
Figure S1: Real space mean field simulations of the slab confirm that assuming constant mean fields is a valid assumption. Mean field result of Eq. (S16) for $t^{\prime}=0.3t$, $U=6t$, $\delta=1$, around half-filling $n=0.54$. (a) The AFM magnetization $m_{z}$ tends to increase slightly towards the surface, whereas the density $n_{z}$ is almost constant. (b) Comparison of the spin splitter effects for layer-dependent mean field results (solid) and using the average values for the mean fields (dashed). (c-f) Comparison of the surface (c-d) and bulk (e-f) spin spectral density (opacity encodes absolute spectral density and the red to blue encodes the spin character) when using the layer-dependent mean field values (c,e) and their mean values (d,f).
III Minimal model for orthorhombic CuMnAs
@@
The bulk polymorph of CuMnAs has orthorhombic space group Pnma (SG #62). Note a tetragonal structure with space group P4/nmm (SG #129) also exists, but is not the focus of this work [52]. Pnma is a non-symmorphic space group generated by three mutually orthogonal mirror symmetries $M_{a}\{\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2}\}$, $M_{b}\{0,\tfrac{1}{2},0\}$ and $M_{c}\{\tfrac{1}{2},0,\tfrac{1}{2}\}$. We take the coordinate system such that $(y,z,x)$ correspond to $(a,b,c)$ axes. All three atoms (Cu,Mn,As) sit on 4c positions [14]. Magnetic moments are located at the Mn atoms in an AFM pattern which preserves $\mathcal{IT}$ symmetry. In addition, in the absence of SOC the symmetry of the magnetic order can be analyzed separately in spin and spatial sectors, and the magnetic order is odd under the mirror $M_{c}$ but even under $M_{a}$ and $M_{b}$. At the surface normal to the $b$ axis, $\mathcal{IT}$ is broken but the combined symmetry $M_{c}\mathcal{T}$ is preserved in the absence of SOC, enforcing a zero total moment [54] and thus qualifying it as an altermagnetic surface.
@@
Figure S2: Tight-binding model for Wyckoff position 4a of Pmna. Filled (empty circles) are sites on even (odd) layers accounted by $\tau$-Pauli matrices, yellow (purple) circles are sites of the A (B) sublattices accounted by $\sigma$-Pauli matrices (a) In-plane and out-of-plane hoppings which preserve all three accidental translations in Eqs. (S23)-(S25). (b) Further neighbor hoppings which keep only $\bm{t}_{a/2,b/2,c/2}$. (c) Further neighbor hoppings which keep only $\bm{t}_{a/2,c/2}$. (d) (110) hopping $t^{\prime}$, which gaps out the nodal line into 4 Dirac cones. (e-h) Spin-dependent hoppings corresponding to the magnetic order with odd $M_{c}\{\tfrac{1}{2},0,\tfrac{1}{2}\}$ symmetry.
The nodal line occurs at $k_{z}=0$, and is protected by the mirror $M_{b}$ which leaves the $k_{z}=0$ plane invariant. The nodal line occurs within a subset of four bands, which come in pairs of opposite $M_{b}$ eigenvalues. Since Mn sits in Wyckoff position 4c, which is invariant under $M_{b}$, a model with a single orbital at 4c necessarily has all four bands which are even or odd under $M_{b}$, which means it is impossible to build a four band model for the nodal line. This means the Wannier centre for such set of four bands must be in another position with multiplicity 4, which can be 4a or 4b. Both positions are not invariant under $M_{b}$ and either can be picked, so we choose 4a. The model is layered in the $b$ direction, with orbitals at $(0,0,0)$, $(\tfrac{1}{2},\tfrac{1}{2},0)$ in the first layer and $(0,0,\tfrac{1}{2})$, $(\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2})$ in the second layer. These sites are shown in Fig. S2(a) and labelled (1-4) in this order. Taking the coordinate system where $(a,b,c)\leftrightarrow(k_{y},k_{z},k_{x})$ and in the basis $(c_{1},c_{2},c_{3},c_{4})$ corresponding to the four sites in Fig. S2(a), we define sublattice $\sigma_{i}$, layer $\tau_{i}$, and spin $s_{i}$ Pauli matrices.
$M_{a}:$ $H(k_{x},k_{y},k_{z})$ $=\sigma_{x}\tau_{x}H(k_{x},-k_{y},k_{z})\sigma_{x}\tau_{x}$ (S18)
$M_{b}:$ $H(k_{x},k_{y},k_{z})$ $=\tau_{x}H(k_{x},k_{y},-k_{z})\tau_{x}$ (S19)
$M_{c}:$ $H(k_{x},k_{y},k_{z})$ $=\sigma_{x}H(-k_{x},k_{y},k_{z})\sigma_{x}$ (S20)
$\mathcal{I}:$ $H(\bm{k})$ $=H(-\bm{k})$ (S21)